# Novel Bounds From The Weak Gravity and Festina Lente Conjectures

### v1

@@ Abstract
We demonstrate that the Weak Gravity Conjecture (WGC) and the Festina-Lente Conjecture (FLC) yield novel bounds on fifth force searches and milli-Charged Particles (mCPs), as well as on the scale of inflation and on the effective Higgs quartic interaction. In particular, we find that combining the FLC with inflation leads to stronger bounds on mCPs than what the simple application of the FLC provides. Furthermore, we have explored the implications of naturalness on both the FLC and WGC, and have found that these conjectures place a lower limit on the charge of a $U(1)$ gauge group.
@@ I Introduction
It is quite known that not all Quantum Field Theories (QFTs) that are consistent at low energies, can have ultraviolet (UV) completions compatible with quantum gravity. QFTs that can be UV completed to include gravity are said to “live” in the landscape, whereas those that cannot are considered to belong to the swampland. The main objective of the Swampland program [1, 2, 3, 4] is to qualify QFTs that can be UV completed to incorporate gravity. This program is supported by a multitude of pieces of evidence from string theory, in addition to other considerations, such as the absence of global symmetries in quantum gravity [5].
A cornerstone of the Swampland program is the Weak Gravity Conjecture (WGC) [6], which roughly states that “gravity is the weakest force”. More precisely, the WGC states that in order for charged black holes in de Sitter (dS) space (known as Reissner-Nordström (RN) black holes) to be able to decay and avoid super-extremality [^1] For an RN black hole with mass $M$, electric charge $Q$ and magnetic charge $P$, the event horizon is located at $r_{\pm}=GM\pm\sqrt{G^{2}M^{2}-G(Q^{2}+P^{2})}$ ($G$ is the gravitational constant). The black hole is called extremal when $GM^{2}=Q^{2}+P^{2}$, whereas for $GM^{2}<Q^{2}+P^{2}$, it is called super-extremal. For a super-extremal black hole, the metic is regular everywhere except at the (naked) singularity found at $r=0$. This violates the Cosmic Censorship Conjecture., for the particle with the smallest mass-to-charge ratio $m/q$ (with $q$ being the charge under any exact local $U(1)$ symmetry) the following inequality
@@
must be satisfied, where $g$ is the gauge coupling and $M_{P}=\sqrt{8\pi/G}=2.43\times 10^{18}$ GeV is the reduced Planck mass. Eq. (1) is called the electric form of the WGC and should also apply to magnetic monopoles, i.e., $m_{\text{mag}}\leq\sqrt{2}g_{\text{mag}}q_{\text{mag}}M_{P}$. On the other hand, $g_{\text{mag}}=1/g_{\text{el}}$, and the mass of the magnetic monopole is at least of the order of the magnetic field it generates, or specifically, $m_{\text{mag}}=\Lambda/g_{\text{el}}^{2}$ on dimensional grounds, in terms of a natural cutoff scale $\Lambda$ where the $U(1)$-based effective field theory (EFT) breaks down. Combining the above statements, one deduces the following constraint
@@
also known as the magnetic form of the WGC. For instance, considering the hypercharge $U(1)_{Y}$ symmetry of the Standard Model (SM), Eq. (2) suggests that the scale of New Physics (NP) is at $\sim 10^{17}$ GeV [6]. The WGC has several interesting phenomenological implications, see for instance [7, 8, 9, 10] (for a comprehensive review of the WGC and its significance, see [11]).
Another proposed conjecture within the Swampland program is the Festina-Lente Conjecture (FLC) [12, 13], which states that mass of every particle charged under a given $U(1)$ gauge group must satisfy the following inequality:
@@
where $H$ is the Hubble scale [^2] The FLC applies only in dS or quasi-dS spacetime. Hence, the Hubble scale in Eq. (3) could be at any cosmological epoch where the spacetime is either dS or quasi-dS, including the universe during inflation., and $V$ is the vacuum energy density.
The main argument of the FLC relies on the evaporation of charged black holes in dS spacetime. More specifically, extremal charged black holes in dS space, where the Schwarzschild radius and the cosmological radius coincide, are called Nariai black holes [14, 15]. These black holes lose their charge via the Hawking radiation, in addition to the Schwinger pair production. The Schwinger pair production rate is given by
@@
where the last step arises from the fact that the typical electric field of a Nariai black hole is of order $E\sim gM_{P}H=g\sqrt{V}$. The basic premise of the FLC is that the Nariai black holes should not discharge too quickly, otherwise they would crunch to a naked singularity, thereby violating the Cosmic Censorship Conjecture. Thus, their discharge via the Schwinger pair production must be exponentially suppressed leading to Eq. (3). Another way to look at Eq. (3) is that a magnetic monopole should lie outside its horizon, i.e. it should not form a black hole, thereby fixing the prefactor in Eq. (3). We see that while the WGC places an upper bound on the mass of a charged particle (specifically, the particle with the smallest $m/q$ ratio in the spectrum), the FLC places a lower bound on the mass of a charged particle (in this case, all charged particles in the spectrum).
@@
The FLC is trivially satisfied in QED of the SM, as can be readily verified by applying Eq. (3) to the electron, the lightest charged particle in the SM. Nevertheless, the FLC carries significant phenomenological implications studied in detail in [13]. Namely, it was established, for instance, that any $SU(N)$ must either confine or undergo spontaneous symmetry breaking. In addition, it was demonstrated that the FLC forbids the Higgs potential from developing a minimum at $\Phi=0$ in the broken phase. In [22], the bounds on millicharged particles (mCPs) arising from mixing between an additional $U(1)$ and the SM $U(1)_{\text{QED}}$ were derived. Furthermore, in [23], the FLC was used to investigate a possible UV vacuum structure of the Higgs field beyond the SM (BSM) and to set bounds on the scale of inflation driven by a BSM Higgs potential.
@@
In this paper, we investigate some further phenomenological implications of the FLC and WGC. In particular, we consider the relevance of the FLC to the fifth force searches, where we show that both the FLC and WGC can set strong bounds on any fifth force, either gauge $U(1)$ or Yukawa-type, and could even exclude the entire parameter space for light force carriers with masses comparable to the Hubble scale. We also investigate the implications of the FLC on the Higgs potential instability and set bounds on the effective quartic coupling of the Higgs boson. In addition, we investigate the implications of the FLC on the scale of inflation. Specifically, we show that combining the FLC with the requirement that the cosmological reheating does not impact the Big Bang Nucleosynthesis (BBN), stringent bounds can be placed on mCPs, which are stronger than those emerging from a simple application of the FLC presented in [22]. We also investigate the implications of naturalness on the FLC combined with the WGC, and show that it sets a lower bound on the charge of a $U(1)$ gauge group.
@@
This paper is organized as follows. In Section II, we present the bounds from both conjectures on fifth force searches. In Section III, we examine the phenomenological implications of the FLC under the assumption that gravity is the weakest force, effectively combining FLC with WGC. In Section IV, we review the magnetic form of the FLC and discuss some of its basic applications to the Higgs instability, inflation, and mCPs. In Section V, we explore the role of the FLC in light of naturalness, and finally we conclude in Section VI.
@@ II The FLC and fifth force searches
$U(r)=-\frac{Gm^{2}}{r}(1+\alpha)e^{-\frac{r}{\lambda}}\,$, (5)
where $\alpha$ parametrizes the strength of the new interaction relative to that of gravity, and the force range $\lambda=1/M$ (not to be confused with the Higgs quartic coupling) in terms of the mass of the force carrier $M$. Thus, it is possible to use both the WGC and FLC to set upper and lower bounds on any additional $U(1)$ interaction. However, there is a caveat when applying the WGC: unlike the FLC, which applies to every particle charged under some $U(1)$ in the spectrum, the WGC only applies to the particle with the smallest $m/q$ ratio. Since for any $U(1)$ such a particle must exist to allow for charged black holes to decay, we can proceed with using the WGC in the context of constraining the fifth force.
@@
$U(r)=-\frac{q^{2}g^{2}}{4\pi}\frac{e^{-\frac{r}{\lambda}}}{r}=\frac{-1}{8\pi M_{P}^{2}}\frac{m^{2}}{r}\Big(\frac{2M_{P}^{2}q^{2}g^{2}}{m^{2}}\Big)e^{-\frac{r}{\lambda}}\,.$ (6)
@@
The latter term suggests that the FLC can provide a meaningful bound only for light particles i.e. with masses not much larger than $H\sim 10^{-33}$ eV. Figure 1 shows the bounds implied by the WGC and by the FLC for some benchmark masses on fifth force searches compared to existing bounds. As the plot shows, both the WGC and the FLC set stringent bounds on the parameter space, and can even exclude it entirely for masses not much larger than the Hubble scale. Notice that the bounds arising from the FLC are independent of the mass of the force carrier $M=1/\lambda$, as can be seen from Eq. (7).
@@
Figure 1: Bounds on the fifth force relative strength $\alpha$ provided by the WGC (blue region) and the FLC (green lines) for a broad range of mediator masses $M$ and the corresponding force ranges $\lambda=1/M$. The FLC bounds are shown for a few benchmark particle masses $m$. The region above each line is excluded for the corresponding benchmark mass. The existing bounds from fifth force searches are shown in gray.
III Gravity as the weakest force
As alluded to in the introduction, the physical content of the WGC is that gravity is weaker than any force associated with a $U(1)$ gauge group. More specifically, consider the gravitational force and a $U(1)$ force mediated via a massless force carrier (like the photon), then the WGC implies that
$F_{\text{grav}}=\frac{m^{2}}{8\pi M_{P}^{2}r^{2}}\leq F_{\text{gauge}}=\frac{q^{2}g^{2}}{4\pi r^{2}}\,$, (8)
which immediately leads to Eq. (1). On the other hand, if we consider the FLC in Eq. (3) and multiply both sides by $(4\pi M_{P}^{2}r^{2})^{-1}$, then rearrange the inequality in the light of Eq. (8), we find that the WGC and FLC give rise to
$\frac{V}{m^{2}M_{P}^{2}}\leq\frac{F_{\text{grav}}}{F_{\text{gauge}}}\leq 1\,$, (9)
This suggests a lower bound on the mass of any particle charged under $U(1)$,
$m\geq\frac{\sqrt{V}}{M_{P}}=\sqrt{3}H\,$, (10)
which is trivially satisfied for the SM QED for $H$ being the Hubble constant at present. As we shall see in Section IV.3, the same results can be derived from the magnetic form of the FLC.
The next question that arises is: What about forces mediated by scalar interactions? There have been several attempts in the literature to extend the WGC to include scalar-mediated interactions, see e.g. [24, 25, 7, 26, 27, 28, 29]. Most of these proposals essentially try extend the physical content of the WGC of gravity being the weakest force, to forces mediated by scalar particles, assuming that the gravitational force should be weaker than that mediated by any (massless) scalar particles. We will consider the leading example presented in [24] here and see the implications of the FLC on such forces.
Consider a (massive) scalar field $\phi$ that is charged under some $U(1)$ with charge $q$, and has a mass that is a function of several massless scalar mediator fields $\varphi_{i}$, i.e., $m=m(\varphi_{i})$. Such particles could arise, for example, in string particles that are functions of the moduli of compactification. Then according to [24], gravitationally-bound states should not be allowed, and the absence of such gravitationally bound states implies that the repulsive $U(1)$ gauge interaction, should be larger than the combined attractive gravitational and scalar-mediated forces. In other words, to avoid gravitationally-bound state, one needs [24]
$q^{2}\geq\frac{m^{2}}{M_{P}^{2}}+g^{ij}(\partial_{\varphi_{i}}m)(\partial_{\varphi_{j}}m)\,$, (11)
where $g^{ij}$ is the metric of the kinetic terms in field space. Now, if $\phi$ is not charged under any $U(1)$, then the absence of any gravitationally bound states requires that the scalar force be both repulsive and stronger than the gravitational force,
$g^{ij}(\partial_{\varphi_{i}}m)(\partial_{\varphi_{j}}m)\geq\frac{m^{2}}{M_{P}^{2}}\,$, (12)
which, for a single scalar mediator, implies that the magnitude of the scalar-mediated force is larger than the magnitude of the gravitational force
$|\partial_{\varphi}m|^{2}\geq\frac{m^{2}}{M_{P}^{2}}\,.$ (13)
When combining this with the mass bound in Eq. (10), one finds that
$|\partial_{\varphi}m|\geq\sqrt{3}\frac{H}{M_{P}}\sim 10^{-60}.$ (14)
This result only applies to charged scalar particles with derivatives-type interactions as explained in [24].
IV The Magnetic FLC and its applications
@@
As explained in detail in [13], the main argument behind the magnetic form of both the WGC and the FLC is that a magnetic monopole should be a fundamental point-like particle and not form a black hole. More specifically, magnetic monopoles are topologically stable objects that are protected by symmetry [^3] In more detail, if a gauge group $G$ spontaneously breaks leaving a residual $U(1)$ gauge group $G\rightarrow H\supset U(1)$, then the stability of magnetic monopoles is guaranteed by a residual $U(1)$, and their topological protection arises if the vacuum of the manifold has a nontrivial second homotopy group $\pi_{2}(G/H)\neq 0$., which implies they cannot collapse to form a black hole. Thus, the exterior of the monopole ($r\gtrsim 1/\Lambda$, where $\Lambda$ is the energy scale at which the EFT breaks down) should be described by a subextermal magnetically charged Reissner-Nordström black hole. This implies, according to [13], that the following condition must be satisfied
@@
where $\tilde{M}=GM$, $\tilde{Q}^{2}n^{2}=\frac{G\pi}{g^{2}}n^{2}$, $l_{4}^{2}=3M_{p}^{2}/V$, $M$ is the mass of the black hole, $g$ is the coupling of the associated $U(1)$ gauge group, $Q_{n}=2\pi n$ is the charge of the monopole, $\ell_{4}$ is the 4D de Sitter radius, $V$ is the vacuum energy, and the solution of the polynomial determines the black hole horizon $r_{-}$ and the cosmic horizon $r_{+}$, i.e., $P(r_{\pm},\tilde{M},\tilde{Q})=0$. In fact, Eq. (15) expresses the condition that $r_{-}\leq\Lambda^{-1}\leq r_{+}$, or alternatively,
$\frac{r_{+}}{l_{4}}=\sqrt{\frac{1}{6}\Big(1-\sqrt{1-12\tilde{Q}^{2}/l_{4}^{2}}\Big)}\leq\frac{1}{\Lambda_{4}}\leq\sqrt{\frac{1}{6}\Big(1+\sqrt{1-12\tilde{Q}^{2}/l_{4}^{2}}\Big)}=\frac{r_{c}^{\text{Nariai}}}{l_{4}}\,.$ (16)
The latter implies the cutoff $\frac{\tilde{Q}^{2}}{l_{4}^{2}}\leq\frac{1}{12}$, which after substituting $\tilde{Q}$ in terms of the coupling $g$ leads to a lower bound,
@@
To get a more usable form of the magnetic FLC, we expand Eq. (16) in powers of small $\tilde{Q}^{2}/l_{4}^{2}$, which gives to the leading order
$\Lambda\lesssim\sqrt{8}gM_{p}\leq\frac{2}{\sqrt{3}}\frac{m^{2}}{H}=\frac{2m^{2}M_{P}}{\sqrt{V}}\,$, (18)

### v2

@@ Abstract
We demonstrate that the Weak Gravity Conjecture (WGC) and the Festina-Lente Conjecture (FLC) yield novel bounds on any hypothetical fifth force and milli-Charged Particles (mCPs), as well as on the scale of inflation and on the effective Higgs quartic interaction. In particular, we find that combining the FLC with inflation leads to stronger bounds on mCPs than what the simple application of the FLC provides. Furthermore, we have explored the implications of naturalness on both the FLC and WGC, and have found that these conjectures place a lower limit on the charge of a $U(1)$ gauge group.
@@ I Introduction
It is well known that not all Quantum Field Theories (QFTs) that are consistent at low energies, can have ultraviolet (UV) completions compatible with quantum gravity. QFTs that can be UV completed to include gravity are said to “live” in the landscape, whereas those that cannot are considered to belong to the swampland. The main objective of the Swampland program [1, 2, 3, 4] is to qualify QFTs that can be UV completed to incorporate gravity. This program is supported by a multitude of pieces of evidence from string theory, in addition to other considerations, such as the absence of global symmetries in quantum gravity [5].
A cornerstone of the Swampland program is the Weak Gravity Conjecture (WGC) [6], which roughly states that “gravity is the weakest force” [^1] For attempted formulations of the WGC for scalars, see [7, 8, 9, 10, 11, 12].. More precisely, the WGC states that in order for charged black holes in anasymptotically-flat spacetime (known as Reissner-Nordström (RN) black holes) to be able to decay and avoid super-extremality [^2] For an RN black hole with mass $M$, electric charge $Q$ and magnetic charge $P$, the event horizon is located at $r_{\pm}=GM\pm\sqrt{G^{2}M^{2}-G(Q^{2}+P^{2})}$ ($G$ is the gravitational constant). The black hole is called extremal when $GM^{2}=Q^{2}+P^{2}$, whereas for $GM^{2}<Q^{2}+P^{2}$, it is called super-extremal. For a super-extremal black hole, the metric is regular everywhere except at the (naked) singularity found at $r=0$. This violates the Cosmic Censorship Conjecture., the particle with the smallest mass-to-charge ratio $m/q$ (with $q$ being the charge under any exact local $U(1)$ symmetry) must obey the following inequality
@@
where $g$ is the gauge coupling and $M_{P}=1/\sqrt{8\pi G}=2.43\times 10^{18}$ GeV is the reduced Planck mass. In (quasi-)de Sitter spacetime the relevant solutions are RN-dS black holes, and the WGC in de Sitter has been analyzed explicitly in [13].In the small-H limit, and for black holes whose Schwarzschild radius is much smaller than the Hubble radius $(r_{s}\ll H^{-1})$, the RN-dS extremality condition reduces to the flat-space form above, which we will use as a local approximation in the quasi-de Sitter regimes considered in this work. Eq. (1) is called the electric form of the WGC and should also apply to magnetic monopoles, i.e., $m_{\text{mag}}\leq\sqrt{2}g_{\text{mag}}q_{\text{mag}}M_{P}$. On the other hand, $g_{\text{mag}}=1/g_{\text{el}}$, and the mass of the magnetic monopole is at least of the order of the magnetic field that it generates, or specifically, $m_{\text{mag}}=\Lambda/g_{\text{el}}^{2}$ on dimensional grounds in terms of a natural cutoff scale $\Lambda$ where the $U(1)$-based effective field theory (EFT) breaks down. Combining the above statements, one deduces the following constraint
@@
also known as the magnetic form of the WGC. For instance, considering the hypercharge $U(1)_{Y}$ symmetry of the Standard Model (SM), Eq. (2) suggests that the scale of New Physics (NP) is at $\sim 10^{17}$ GeV [6]. The WGC has several interesting phenomenological implications, see for instance [14, 15, 16, 17], and see [18]for a comprehensive review of the WGC and its significance.
Another proposed conjecture within the Swampland program is the Festina-Lente Conjecture (FLC) [19, 20], which states that the mass of every particle charged under a given $U(1)$ gauge group must satisfy the following inequality:
@@
where $H$ is the Hubble scale [^3] The FLC applies only in dS or quasi-dS space. Hence, the Hubble scale in Eq. (3) could be at any cosmological epoch where the space is either dS or quasi-dS, including the universe during inflation., and $V$ is the vacuum energy density.
The main argument of the FLC relies on the evaporation of charged black holes in dS spacetime. More specifically, extremal charged black holes in dS space, where the Schwarzschild radius and the cosmological radius coincide, are called Nariai black holes [21, 22]. These black holes lose their charge via the Hawking radiation, in addition to the Schwinger pair production. The Schwinger pair production rate is given by:
@@
where the last step arises from the fact that the typical electric field of a Nariai black hole is of order $E\sim gM_{P}H=g\sqrt{V}$. The basic premise of the FLC is that the Nariai black holes should not discharge too quickly, otherwise they would crunch to a naked singularity, thereby violating the Cosmic Censorship Conjecture. Thus, their discharge via the Schwinger pair production must be exponentially suppressed, leading to Eq. (3). Another way to look at Eq. (3) is that a magnetic monopole should lie outside its horizon, i.e. it should not form a black hole, thereby fixing the prefactor in Eq. (3). We see that while the WGC places an upper bound on the mass of a charged particle (specifically, the particle with the smallest $m/q$ ratio in the spectrum), the FLC places a lower bound on the mass of a charged particle (in this case, all charged particles in the spectrum).
@@
The FLC is trivially satisfied in QED of the SM, as can be readily verified by applying Eq. (3) to the electron, the lightest charged particle in the SM. Nevertheless, the FLC carries significant phenomenological implications studied in detail in [20]. Namely, it was established, for instance, that any $SU(N)$ must either confine or undergo spontaneous symmetry breaking. In addition, it was demonstrated that the FLC forbids the Higgs potential from developing a minimum at $\Phi=0$ in the broken phase. In [23], the bounds on millicharged particles (mCPs) arising from the mixing between an additional $U(1)$ and the SM $U(1)_{\text{QED}}$ were derived. Furthermore, in [24], the FLC was used to investigate a possible UV vacuum structure of the Higgs field beyond the SM (BSM) and to set bounds on the scale of inflation driven by a BSM Higgs potential.
@@
In this paper, we investigate some further phenomenological implications of the FLC and the WGC. In particular, we consider the relevance of the FLC to the fifth force searches, where we show that both the FLC and WGC can set strong bounds on any fifth force, either gauge $U(1)$ or Yukawa-type. Applied to hypothetical long-range forces, the combined WGC/FLC logic implies a strong parametric constraint; in the extreme IR (mediator masses $(m\sim H))$ this becomes particularly restrictive. We include this limit only to illustrate the scaling with (H), and we do not interpret it as a robust present-day phenomenological exclusion. We also investigate the implications of the FLC on the Higgs potential instability and set bounds on the effective quartic coupling of the Higgs boson. In addition, we investigate the implications of the FLC on the scale of inflation. Specifically, we show that by combining the FLC with the requirement that the cosmological reheating does not impact the Big Bang Nucleosynthesis (BBN), stringent bounds can be placed on mCPs, which are stronger than those emerging from a simple application of the FLC presented in [23]. We also investigate the implications of naturalness on the FLC combined with the WGC, and show that it sets a lower bound on the charge of a $U(1)$ gauge group.
@@
This paper is organized as follows. In Section II, we present the bounds from both conjectures on hypothetical fifth forces. In Section III, we review the magnetic form of the FLC and discuss some of its basic applications to the Higgs instability, inflation, and mCPs. In Section IV, we explore the role of the FLC in light of naturalness, and finally, we conclude in Section V .
@@ II The FLC and fifth force searches
$U(r)=-\frac{Gm^{2}}{r}(1+\alpha e^{-\frac{r}{\lambda}})\,$, (5)
where $\alpha$ parametrizes the strength of the new interaction relative to that of gravity, and the force range $\lambda=1/M$ (not to be confused with the Higgs quartic coupling) in terms of the mass of the force carrier $M$. Thus, it is possible to use both the WGC and the FLC to set upper and lower bounds on any additional $U(1)$ interaction  [^4] We need to point out here that both the FLC and the WGC are based on the discharge and decay of black holes, which were derived for long-range forces (massless U(1) gauge bosons), however, here we are extending the bounds to massive force carriers and not limiting ourselves to the massless regime. So, we are assuming that the decay of BHs via the Hawking radiation, and their discharge via the Schwinger pair production, are not significantly affected by the force carrier being massive instead of massless, as long the mass is not too large.. However, there is a caveat when applying the WGC: unlike the FLC, which applies to every particle charged under some $U(1)$ in the spectrum, the WGC only applies to the particle with the smallest $m/q$ ratio. Since for any $U(1)$ such a particle must exist to allow for charged black holes to decay, we can proceed with using the WGC in the context of constraining the fifth force.
@@
$U(r)=-\frac{q^{2}g^{2}}{4\pi}\frac{e^{-\frac{r}{\lambda}}}{r}=\frac{-1}{8\pi M_{P}^{2}}\frac{m^{2}}{r}\Big(\frac{2q^{2}g^{2}M_{P}^{2}}{m^{2}}\Big)e^{-\frac{r}{\lambda}}\,.$ (6)
@@
The latter term suggests that the FLC can provide a meaningful bound only for light particles i.e. with masses not much larger than $H\sim 10^{-33}$ eV. Figure II shows the bounds implied by the WGC and by the FLC for some benchmark masses on a hypothetical fifth force compared to existing bounds. As the plot shows, both the WGC and the FLC set stringent bounds on the parameter space, and can even disfavor it entirely for masses not much larger than the Hubble scale. Notice that the bounds arising from the FLC are independent of the mass of the force carrier $M=1/\lambda$, as can be seen from Eq. (7). We should point out however, that the bounds obtained from the FLC are based on hypothetical masses of the force carriers and only serve as an illustration. In reality, the mass of the force carrier of a potential fifth force would most probably be much larger than $H$, which would lead to a much weaker bound.
@@
Bounds on the fifth force relative strength $\alpha$ provided by the WGC (blue region) and the FLC (green lines) for a broad range of mediator masses $M$ and the corresponding force ranges $\lambda=1/M$. The FLC bounds are shown for a few benchmark particle masses $m$. The region above each line is disfavored for the corresponding benchmark mass. Regions above the curve would be in tension with the FLC for the chosen benchmark. The existing bounds from fifth force searches are shown in gray.
III The Magnetic FLC and its applications
@@
As explained in detail in [20], the main argument behind the magnetic form of both the WGC and the FLC is that a magnetic monopole should be a fundamental point-like particle and not form a black hole. More specifically, magnetic monopoles are topologically stable objects that are protected by symmetry [^5] In more detail, if a gauge group $G$ spontaneously breaks leaving a residual $U(1)$ gauge group $G\rightarrow H\supset U(1)$, then the stability of magnetic monopoles is guaranteed by the residual $U(1)$, and their topological protection arises if the vacuum of the manifold has a nontrivial second homotopy group $\pi_{2}(G/H)\neq 0$., which implies they cannot collapse to form a black hole. Thus, the exterior of the monopole ($r\gtrsim 1/\Lambda$, where $\Lambda$ is the energy scale at which the EFT breaks down) should be described by a subextremal magnetically charged Reissner-Nordström black hole. This implies, according to [20], that the following condition must be satisfied
@@
where $\tilde{M}=GM$, $\tilde{Q}^{2}n^{2}=\frac{G\pi}{g^{2}}n^{2}$, $\ell_{4}^{2}=3M_{p}^{2}/V$ , $M$ is the mass of the black hole, $g$ is the coupling of the associated $U(1)$ gauge group, $Q_{n}=2\pi n$ is the charge of the monopole, $\ell_{4}$ is the 4D de Sitter radius, $V$ is the vacuum energy, and the solution of the polynomial determines the black hole horizon $r_{-}$ and the cosmic horizon $r_{+}$, i.e., $P(r_{\pm},\tilde{M},\tilde{Q})=0$. In fact, Eq. (8) expresses the condition that $r_{-}\leq\Lambda^{-1}\leq r_{+}$, or alternatively,
$\frac{r_{+}}{\ell_{4}}=\sqrt{\frac{1}{6}\Big(1-\sqrt{1-12\tilde{Q}^{2}/\ell_{4}^{2}}\Big)}\leq\frac{1}{\Lambda_{4}\ell_{4}}\leq\sqrt{\frac{1}{6}\Big(1+\sqrt{1-12\tilde{Q}^{2}/\ell_{4}^{2}}\Big)}=\frac{r_{c}^{\text{Nariai}}}{\ell_{4}}\,.$ (9)
The latter implies the cutoff $\frac{\tilde{Q}^{2}}{\ell_{4}^{2}}\leq\frac{1}{12}$, which after substituting $\tilde{Q}$ in terms of the coupling $g$ leads to a lower bound,
@@
To get a more usable form of the magnetic FLC, we expand Eq. (9) in powers of small $\tilde{Q}^{2}/\ell_{4}^{2}$, which gives to the leading order
$\Lambda\lesssim\sqrt{8}gM_{p}\leq<\frac{2}{\sqrt{3}}\frac{m^{2}}{H}=\frac{2m^{2}M_{P}}{\sqrt{V}}\,$, (11)